# Weiqi Wang

### v1

@@ 5.3 Analysis
5.3.1 Improving Metaphysical Reasoning via Transferring from Conceptualization Taxonomy
Improving the performance of LLMs on Mars requires extensive fine-tuning on large-scale human-annotated data, making it non-trivial. Given that approximately 80% of action changes are executed by modifying a component along with its abstracted concepts, we first study whether exposing LLMs to more conceptualizations and abstract knowledge can enhance their metaphysical reasoning capabilities. For this purpose, we select CANDLE [64] as the knowledge source, which is an automatically constructed knowledge base containing 382K conceptualizations of events and abstract inferential knowledge. We first convert event-conceptualization pairs into the task format of metaphysical event discrimination and reformat commonsense inferential knowledge to align with the objectives of the metaphysical inference and transition discrimination tasks. More details are in Appendix D.2.
Three backbone models are then fine-tuned separately on CANDLE and Mars. Another group is pre-trained on CANDLE before being fine-tuned on Mars. All models are then evaluated on the testing set of Mars, with the results reported in Table 4. From the results, a significant improvement is observed across all tasks when the models are sequentially fine-tuned on CANDLE and Mars, compared to solely fine-tuning on CANDLE or Mars. These findings indicate that the transfer of conceptualizations and abstract knowledge from CANDLE effectively enhances the performance of LMs in metaphysical reasoning tasks. Since CANDLE is constructed by distilling from ChatGPT without human labor, this opens up a scalable and cost-efficient approach to improving the metaphysical reasoning capabilities of LLMs.
5.3.2 Impact of Component Type on LLM’s Performance
We then analyze the performance of LLMs on each component type to understand the reasons for their subpar performance. We select LLaMa-3-8B as the representative model and compare its accuracy on each component type when fine-tuned on Mars and CANDLE + Mars. The results are illustrated in Figure 5. We observe that while pre-training the model on CANDLE consistently enhances performance, LLaMa3 still struggles when reasoning with changes in spatial quantifiers, temporal quantifiers, and numerical properties. This is in line with recent studies that demonstrate weaknesses in temporal and numerical reasoning for LLMs [57, 55]. Another possible reason is that since CANDLE only contains conceptualizations for subjects, verbs, objects, and sub-events in social events, pre-training models on it cannot provide benefits for the aforementioned aspects of change. Moreover, we only observe limited improvement for the metaphysical event discrimination task. This sheds light on future works that could focus on how to further enhance LLM’s metaphysical reasoning capabilities in these weaker dimensions.
@@ 5.3.3 Error Analysis of GPT-Series Models
Finally, we select GPT4 as a representative model and conduct a manual analysis to identify the causes of errors by categorizing the mistakes found in their COT responses. We sample 150 COT responses from each task, all of which result in inconsistent results compared to human annotated labels and present our classifications of these errors as follows:
  • Hallucinations: 41.7% of errors are caused by factual or metaphysical hallucinations by GPT4, where it creates a context that accommodates changes in actions and inferences that are not mentioned in the original text [39, 73]. For instance, in the event “The poet enjoys writing poems about western festivals,” GPT4 incorrectly interprets the poet as Du Fu, an ancient Chinese poet. This leads to a conflict when reasoning about his life and the subsequent inference “He was famous in the west,” resulting in faulty reasoning.
  • Confusion between Concepts and Hypernyms: 36.3% errors are attributed to GPT4’s tendency to perceive abstract components within changed actions as hypernyms that fulfill the change, without considering all potential entities within the original concept. For instance, in a modified event, “He jumps down from very high altitude and lands peacefully,” GPT4 interprets very high altitude as a diving platform, deeming it plausible. However, this concept could also encompass high buildings, which would not be suitable for the event.
  • Internal Conflict: 17.7% errors are attributed to internal conflicts within GPT4’s reasoning rationales, as well as inconsistencies between the binary predictions made and the corresponding reasoning rationales.
  • Annotation Error: 4.3% errors are erroneously identified due to incorrect labels, potentially caused by spamming or a misunderstanding of the task by human annotators.
@@ 6 Conclusions
In conclusion, this paper proposes Metaphysical Reasoning to delineate the process of reasoning with changes in distribution and construct [Uncaptioned image]Mars as the associated evaluation benchmark in a non-trivial manner. Our extensive experiments show the difficulty of our proposed task, which cannot be easily addressed through advanced prompting and fine-tuning. Further analysis reveals why LMs encounter difficulties with metaphysical reasoning tasks and suggests a potential avenue for improvement. We hope that our work illuminates the path towards achieving conscious processing in LLMs via System II reasoning through effective comprehension of changes in distribution.
Acknowledgments and Disclosure of Funding
The authors of this paper thank Tianqing Fang, Wenxuan Ding, Wei Fan, Jiaxin Bai, and the anonymous reviewers for their constructive feedbacks. The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
Appendix A Limitations and Future Works
@@
Though we consider our work to be a fundamental step towards understanding the capabilities of LMs in reasoning with changes in distribution, several limitations still exist that cannot be covered within this single work. Here, we discuss some important limitations that future works can address:
@@
  • Expand current formulation of changes to include more types of components. In our formulation, we model changes within an action by altering a component within that action to describe the change that has occurred. This approach provides a structured ontology for formulating the various changes that can potentially occur within an event. In our work, we focus on seven primary types of changes, covering the subject, verb, object, spatial quantifier, temporal quantifier, numerical properties, and sub-events of the event. While these seven types encompass most of the potential changes, there are other components within an event that can be modified, such as adjectives, adverbs, and prepositional phrases. Despite this, our benchmark curation pipeline is flexible regarding the types of components it can handle. By incorporating human-authored examples, ChatGPT can perform automatic extraction, thereby easily covering a broader range of component types. Consequently, future research can extend our benchmark to include these additional components.
  • Enabling simultaneous modification of multiple components. For simplicity, we consider only one change occurring per event in each data entry. However, it is also possible for multiple changes to occur simultaneously, thereby modifying more than one component. This, however, could lead to a significantly larger dataset, rendering it impractical to construct a benchmark with human annotation.
  • Implement more detailed changes through instantiation. In our work, we use abstractions to model changes in an event with a hierarchical ontology. The higher the level of abstraction, the more changes it can represent. To reflect these abstracted changes back to their original level in the event, future studies could instantiate them to obtain more detailed changes.
@@
  • Reliance of ChatGPT on benchmark curation. Our data construction process relies significantly on ChatGPT, an expensive and proprietary language model used for data collection, as well as human annotation for data verification. Future research could consider utilizing robust open-source language models [53] and general statement plausibility estimators [40] to replace these methods. This could potentially reduce costs and construct larger benchmark.
Appendix B Ethics Statement
B.1 Offensive Content Elimination
@@
Our benchmark curation pipeline, which involves generating content with ChatGPT, necessitates stringent measures to ensure the absence of offensive content in both the prompts and the generated responses. For this purpose, we apply two strategies to eliminate offensive content. First, we use the highest level of Azure AI Content Safety Filter to filter out any content that contains personal privacy, promotes violence, racial discrimination, hate speech, sexual content, or self-harm. If any such unsafe content is detected in the prompts or generated responses, it automatically triggers a system failure, which prevents the inclusion of such data in our dataset. Second, we manually inspect a random sample of 500 data entries from three tasks in [Uncaptioned image]Mars for offensive content. Based on our annotations, we have not detected any offensive content. We thus believe that our dataset is safe.
B.2 Licenses
@@
We will share our code, dataset, and models under the MIT license, thereby granting other researchers free access to our assets for research purposes. If there are any issues with rights violations or anything related to the data license, we will take full responsibility. Other datasets used in this paper, including Wikitext and Bookcorpus, are shared under the CC-SA license, permitting us to use them for research. As for language models, we access all open-source LMs via the Huggingface Hub [70]. All associated licenses permit user access for research purposes, and we have agreed to follow all terms of use.
B.3 Annotations
@@
We conduct large scale human annotations on the Amazon Mechanical Turk (AMT) platform. The annotators are paid on average at an hourly rate of 19 USD, which is comparable to the minimum wages in the US. The selection of these annotators is solely based on their performance on the evaluation set, and we do not collect any personal information about the participants from AMT. For expert verifications, we have secured IRB approval and support from our institution’s department, which allows us to invite expert graduate students to validate the quality of our data. They all agree to participate voluntarily without being compensated. We have made concerted efforts to eliminate offensive content, thereby ensuring that no annotators are offended.
B.4 Societal Impact
Given that [Uncaptioned image]Mars solely contains events, inferences, and transitions that are either metaphysical or plausible, and considering that we have meticulously removed any offensive content, we believe that this paper will not have any negative societal impact.
Appendix C Prompts for Generations and Evaluations with LLMs
C.1 [Uncaptioned image]Mars Benchmark Curation
@@ Appendix C Prompts for Generations and Evaluations with LLMs
We first present our prompts used in each step for sequentially instructing ChatGPT to generate candidate data for [Uncaptioned image]Mars.
C.1.1 Text Decomposition and Event Component Extraction
@@
C.1.2 Component Abstraction and Variation
@@
C.1.3 Inference Generation
@@
Note that leveraging LLM to perform contextualized abstraction [64, 75] has been shown to result in larger coverage and stronger downstream benefits compared to previous conceptualization methods [23], such as retrieving from a pre-defined concept taxonomy or human annotation. Our method is justifiable and enables large-scale benchmark construction.
C.1.4 Metaphysical Transition Generation

### v2

@@ 5 Evaluations and Analysis
5.2 Main Evaluations on [Uncaptioned image]Mars
5.2.1 Task Setup and Model Selections
We then experiment with a selection of (L)LMs to investigate their performances on our curated Mars benchmark. Accuracy, AUC, and Macro-F1 scores are used as evaluation metrics.
The evaluation of different models are categorized into three types: (1) Zero-shot: We first evaluate several (L)LMs in a zero-shot manner. For small-sized Pre-Trained Language Models (PTLMs), we evaluate DeBERTa-v3 He et al. (2023a), GPT2 Radford et al. (2019), CAR Wang et al. (2023a), CANDLE Wang et al. (2024b), and VERA Liu et al. (2023a), following the design of zero-shot question answering Ma et al. (2021). For LLMs, we evaluate LLaMa2, LLaMa3, LLaMa3.1 Touvron et al. (2023a, b); Dubey et al. (2024), Gemma Mesnard et al. (2024), Falcon Almazrouei et al. (2023), and Mistral Jiang et al. (2023) using direct zero-shot prompting Qin et al. (2023). (2) Finetuning: We then assess the performance of (L)LMs when fine-tuned on the training set of Mars. For PTLMs, we fine-tune DeBERTa, GPT2-xl, and VERA. For LLMs, we fine-tune LLaMa2, LLaMa3, Gemma, and Mistral using LoRA Hu et al. (2022). (3) LLM API: Finally, we evaluate the performance of GPT-4 OpenAI (2023) and GPT-4o-mini OpenAI (2024), which represent proprietary LLMs, under zero-shot, five-shots, Chain-of-Thought prompting (COT; Wei et al., 2022), and Self-Consistent COT (SC-COT; Wang et al., 2023c) settings. For LLaMa3.1-70B and GPT-4o-mini, we also test their performances with RAG Gao et al. (2023), Multi-agent Calibration Yang et al. (2024), and Self Reflection Pan et al. (2024). Please find implementation details in Appendix C, multi-task fine-tuning experiments in Appendix E.1, and few-shot fine-tuning experiments in Appendix E.2.
5.2.2 Results and Analysis
Evaluation results are reported in Table 2. From the results, we observe that: (1) Most models exhibit subpar performance under the zero-shot setting. Among PTLMs, only VERA delivers acceptable results across all three tasks, while the rest significantly underperform. Though models fine-tuned on commonsense knowledge and conceptualizations, such as CAR and CANDLE, show some improvement compared to their DeBERTa-v3-Large backbone, these performances are still unsatisfactory, even falling below the level of majority voting. For LLMs, improving training paradigms and increasing the number of parameters can indeed help achieve better performance. Nevertheless, all models perform poorly across all tasks in Mars, emphasizing the difficulty of our tasks. (2) Fine-tuning only offers limited benefits. With fine-tuning, all models improve significantly. For example, DeBERTa-Large’s accuracy increases by 16.18%, 21.84%, and 22.2% on three tasks, respectively. However, the best results for all tasks are still capped at around 74%, indicating a shared difficulty and significant room for future enhancements. One potential reason for this is that, since we split the data according to the source of text in Wikitext and BookCorpus, the distribution between different splits may differ significantly, as the domain and topics could be diverse from each other. We also discuss the reasons for PTLMs’ strong performance compared to LLMs after fine-tuning in Appendix E.3. (3) The GPT series models underperform compared to other LLMs, and COT does not consistently aid performance. Surprisingly, GPT series models fall short when compared to open LLMs, such as LLaMa-3-70B. One possible explanation is that negative examples in Mars are sourced from ChatGPT’s generation and are obtained via post-human annotation. This makes it challenging to discriminate as these negative examples contradict GPT’s internal knowledge. Advanced prompting methods only offer limited improvement in performances.
@@ 5.3 Analysis
5.3.1 Transferring from Conceptualization
Improving the performance of LLMs on Mars requires extensive fine-tuning on large-scale human-annotated data, making it non-trivial. Since we observe that approximately 80% of action changes are executed by modifying a component along with its abstracted concepts (see Table 5), we first study whether exposing LLMs to more conceptualizations and abstract knowledge can enhance their metaphysical reasoning capabilities. For this purpose, we select CANDLE Wang et al. (2024b) as the knowledge source, which is an automatically constructed knowledge base containing 382K conceptualizations of events and abstract inferential knowledge. We first convert event-conceptualization pairs into the task format of metaphysical event discrimination and reformat commonsense inferential knowledge to align with the objectives of the metaphysical inference and transition discrimination tasks. More details are in Appendix C.2.
Three backbone models are then fine-tuned separately on CANDLE and Mars. Another group is sequentially fine-tuned on CANDLE and then on Mars. All models are then evaluated on the testing set of Mars, with the results reported in Table 3. From the results, a significant improvement is observed across all tasks when the models are sequentially fine-tuned on CANDLE and Mars, compared to solely fine-tuning on CANDLE or Mars.
These findings indicate that the transfer of conceptualizations and abstract knowledge from CANDLE effectively enhances the performance of LMs in metaphysical reasoning tasks. Since CANDLE is constructed by distilling from an LLM without human labor, this opens up a scalable and cost-efficient approach to improving the metaphysical reasoning capabilities of LLMs.
5.3.2 Impact of Component Types
We then analyze the performance of LLMs on each component type to understand the reasons for their subpar performance. We select LLaMa-3-8B as the representative model and compare its accuracy on each component type when fine-tuned on Mars and CANDLE + Mars. The results are illustrated in Figure 4. We observe that while pre-training the model on CANDLE consistently enhances performance, LLaMa3 still struggles when reasoning with changes in spatial quantifiers, temporal quantifiers, and numerical properties. This is in line with recent studies that demonstrate weaknesses in temporal and numerical reasoning for LLMs Tan et al. (2023); Shi et al. (2023). Another possible reason is that since CANDLE only contains conceptualizations for subjects, verbs, objects, and sub-events in social events, pre-training models on it cannot provide benefits for the aforementioned aspects of change. Moreover, we only observe limited improvement for the metaphysical event discrimination task. Future works could focus on how to further enhance LLM’s metaphysical reasoning capabilities in these weaker dimensions.
@@ 5.3.3 Error Analysis of GPT-Series Models
Finally, we select GPT4 as a representative model and conduct a manual analysis to identify the causes of errors by categorizing the mistakes found in their COT responses. We sample 150 COT responses from each task, all of which result in inconsistent results compared to human annotated labels and present our classifications of these errors as follows: (1) Hallucinations: 41.7% of errors are caused by factual or metaphysical hallucinations by GPT4, where it creates a context that accommodates changes in actions and inferences that are not mentioned in the original text. For instance, in the event “The poet enjoys writing poems about western festivals,” GPT4 incorrectly interprets the poet as Du Fu. This leads to a conflict when reasoning about his life and the subsequent inference “He was famous in the west,” resulting in faulty reasoning. (2) Confusion between Concepts and Hypernyms: 36.3% errors are attributed to GPT4’s tendency to perceive abstract components within changed actions as hypernyms that fulfill the change, without considering all potential entities within the original concept. For instance, in a modified event, “He jumps down from very high altitude and lands peacefully,” GPT4 interprets very high altitude as a diving platform, deeming it plausible. However, this concept could also encompass high buildings, which would not be suitable for the event. (3) Internal Conflict: 17.7% errors are attributed to internal conflicts within GPT4’s reasoning rationales, as well as inconsistencies between the binary predictions made and the corresponding reasoning rationales. (4) Annotation Error: 4.3% errors are erroneously identified due to incorrect labels, potentially caused by spamming or a misunderstanding of the task by human annotators.
@@ 6 Conclusions
In conclusion, this paper proposes Metaphysical Reasoning to delineate the process of reasoning with changes in distribution and construct [Uncaptioned image]Mars as the associated evaluation benchmark in a non-trivial manner. Our experiments show the challenge of our task, which advanced prompting and fine-tuning can’t easily solve. Analysis reveals why LMs struggle with metaphysical reasoning and suggests a possible improvement. We hope to illuminate the path toward achieving conscious processing in LLMs through System II reasoning by effectively comprehending changes in distribution.
Limitations
@@
Though we consider our work to be a fundamental step towards understanding the capabilities of LMs in reasoning with changes in distribution, we do acknowledge that several limitations still exist that just cannot be covered within one single work. Here, we discuss some important limitations that future works can address:
@@
(1) Include more types of changes in our current formulation. In our work, we primarily focus on seven types of changes, covering the subject, verb, object, spatial quantifier, temporal quantifier, numerical properties, and sub-events of the event. While these seven types encompass most of the potential changes, there are other uncovered components within an event that can be impacted by changes, such as adjectives, adverbs, and prepositional phrases. Nevertheless, our flexible and automated benchmark curation pipeline, empowered by an LLM, allows for future research to extend the benchmark to cover a broader range of component types.
@@
(2) Reliance of LLM on benchmark curation. Our data construction process relies significantly on ChatGPT, an expensive and proprietary language model used for data collection, as well as human annotation for data verification. In Appendix B.3, we discussed the feasibility of leveraging open-sourced LLM as a replacement to ChatGPT to reduce cost and promote reproducibility. Future research could also consider utilizing robust open-source language models Reid et al. (2024) and general statement plausibility estimators Liu et al. (2023a) to replace these methods.
(3) Solution and Downstream Applications of Metaphysical Reasoning. While this paper establishes a comprehensive evaluation benchmark for metaphysical reasoning, we leave the exploration of a practical solution to aid LLMs in solving metaphysical reasoning tasks, as well as the potential benefits of utilizing metaphysical reasoning for downstream tasks into future works. These tasks may include planning Yuan et al. (2023); Ouyang and Li (2023) or reasoning with changes He et al. (2023b).
Ethics Statement
@@
Offensive Content Elimination. Our benchmark curation pipeline, which involves generating content with ChatGPT, necessitates stringent measures to ensure the absence of offensive content in both the prompts and the generated responses. For this purpose, we apply two strategies to eliminate offensive content. First, we use the highest level of Azure AI Content Safety Filter to filter out any content that contains personal privacy, promotes violence, racial discrimination, hate speech, sexual content, or self-harm. If any such unsafe content is detected in the prompts or generated responses, it automatically triggers a system failure, which prevents the inclusion of such data in our dataset. Second, we manually inspect a random sample of 500 data entries from three tasks in [Uncaptioned image]Mars for offensive content. Based on our annotations, we have not detected any offensive content. We thus believe that our dataset is safe and will not yield any negative societal impact.
@@
Licenses. We will share our code and models under the MIT license, thereby granting other researchers free access to our assets for research purposes. Other datasets used in this paper, including Wikitext and Bookcorpus, are shared under the CC-SA license, permitting us to use them for research. As for language models, we access all open-source LMs via the Huggingface Hub Wolf et al. (2020). All associated licenses permit user access for research purposes, and we have agreed and committed to follow all terms of use.
@@
Annotations. We conduct large scale human annotations on the Amazon Mechanical Turk (AMT) platform. We invite annotation workers from the US, Europe, and India due to their proficiency in English. The annotators are paid on average at an hourly rate of 19 USD, which is comparable to the minimum wages in the US. The selection of these annotators is solely based on their performance on the evaluation set, and we do not collect any personal information about the participants from AMT. For expert verifications, we have secured IRB approval and support from our institution’s department, which allows us to invite expert graduate students to validate the quality of our data. They all agree to participate voluntarily without being compensated. We have made concerted efforts to eliminate offensive content, thereby ensuring that no annotators are offended.
Acknowledgements
We thank the anonymous reviewers and the area chair for their constructive comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from the ITC of Hong Kong, SAR, China, as well as the AoE (AoE/E-601/24-N), the RIF (R6021-20), and the GRF (16205322) from the RGC of Hong Kong, SAR, China.
Appendices
Appendix A Differentiation from Philosophical Metaphysics and Counterfactual Reasoning
In this work, we use the term “metaphysical” to describe a specific mode of reasoning that deals with highly improbable or abstract scenarios, distinct from both its traditional philosophical meaning and the concept of counterfactual reasoning. Philosophically, “metaphysics” refers to the study of the fundamental nature of reality, encompassing questions about existence, causality, and the nature of being Aristotle and Aristotle (1933); Bergson (1999). While this classical usage involves conceptual analysis and abstract thought, our focus diverges significantly. We adopt “metaphysical” to signify reasoning that examines transitions between plausible and highly improbable states, emphasizing the logical structure and abstracted nature of these transitions rather than ontological or existential inquiries.
This distinction is important because our framework does not engage with the philosophical debates about the nature of reality or existence. Instead, it concentrates on how LLMs process and adapt to scenarios that are rare or abstract yet logically consistent. For example, while metaphysical reasoning in our context might involve reasoning about a scenario where “a civilization survives for 100,000 years,” it does not explore the metaphysical nature of time, existence, or causality in a philosophical sense.
Furthermore, our concept of metaphysical reasoning is distinct from counterfactual reasoning. Counterfactual reasoning involves evaluating “what if” scenarios that diverge from known realities but remain bounded by plausible causal relationships Li et al. (2023); Hua et al. (2024). For example, a counterfactual might consider, “What if Caesar had lost the battle of Pharsalus?”–a scenario grounded in historical plausibility. In contrast, metaphysical reasoning in our framework extends beyond plausibility to explore scenarios that are structurally coherent but unlikely or abstract, such as “What if Caesar ruled for a millennium?” Here, the focus is not on causal plausibility but on the ability to evaluate transitions to rare, abstract, or highly improbable states.
This differentiation between “metaphysical” in our framework, metaphysics in philosophy, and counterfactual reasoning underscores the novel challenges our benchmarks aim to address. By pushing LLMs to reason about transitions into abstract or improbable scenarios, we aim to probe and enhance their capabilities for adaptive, out-of-distribution reasoning – a necessary step toward achieving generalizable System II reasoning.
Appendix B Mars Benchmark Curation Details
B.1 [Uncaptioned image]Mars Benchmark Curation
@@ Appendix B Mars Benchmark Curation Details
An overview of our benchmark construction pipeline is shown in Figure 5. We first present our prompts used in each step for sequentially instructing ChatGPT to generate candidate data for [Uncaptioned image]Mars Wang et al. (2024a).
B.1.1 Text Decomposition and Event Component Extraction
@@
B.1.2 Component Abstraction and Variation
@@
Note that leveraging LLM to perform contextualized abstraction Wang et al. (2024b); Yu et al. (2023) has been shown to result in better quality, larger coverage, and stronger downstream benefits compared to previous conceptualization methods He et al. (2024); Wang et al. (2023b, 2024c), such as retrieving from a pre-defined concept taxonomy or human annotation. Our knowledge distillation-based method is justifiable and enables large-scale benchmark construction.
B.1.3 Inference Generation
@@
Model Task 1 Plaus. Expert. Task 2 Plaus. Expert. Task 3 Plaus. Expert.
ChatGPT 60.98 92.0 58.56 96.5 50.25 93.5
Meta-LLaMa-3.1-405B 62.2 93.2 57.0 95.8 51.0 94.6
GPT-4o 64.6 94.8 59.2 98.4 53.4 96.0
Table 4: Annotation results of evaluation data curated with different LLMs as backbones. Plaus. refers to plausible event/inference/transition rate and Expert. refers to ratio of data accepted by expert annotators.
B.1.4 Metaphysical Transition Generation